OpenStudy (anonymous):

Write in standard form. 11^2 x 3^3

OpenStudy (anonymous):

Well, what's 11 * 11?

OpenStudy (anonymous):

121

OpenStudy (anonymous):

3 * 3 * 3 = ?

OpenStudy (anonymous):

\[x^2 = x * x \]

OpenStudy (anonymous):

27

OpenStudy (anonymous):

\[x^3 = x * x * x \]

OpenStudy (anonymous):

Well, actually I don't think you need to multiply for this. The answer is asking for standard form. So, one has ^2 and one has ^3. ^3 is bigger than ^2 so, you just switch them around. Understand that? If you do, show me what you get.

OpenStudy (anonymous):

Do you have options?

OpenStudy (anonymous):

i dont understand.

OpenStudy (anonymous):

Standard form is biggest to lowest amount. Example: You have 5x and 7x Standard form of it is: 7x + 5x

OpenStudy (anonymous):

Do you understand?.-.

OpenStudy (anonymous):

okay i got that

OpenStudy (anonymous):

So, input standard form into your probably where bigger and lowest amount account for.

OpenStudy (anonymous):

11 x 11x 3 x 3 x 3 ?

OpenStudy (anonymous):

Not quite. You don't have to multiply at all. You just need to arrange the amounts from biggest to lowest.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So what do you get?

OpenStudy (anonymous):

idk :P

OpenStudy (anonymous):

Well if you read my big comment above, I said ^3 is bigger than ^2 :P

OpenStudy (anonymous):

yeah .. :P

OpenStudy (anonymous):

3^3 x 11^2 it's that easy :P I told you all you had to do was switch them. Needa read what I say :b

OpenStudy (anonymous):

okay thx

OpenStudy (anonymous):

Welcome.